I am taking and entrance exam to get into a trade school Im having problems with these type of problems.Can you explain how to come up with the answer to some these questions namely no.2? Whatever help you can offer me would be greatly appreciated.
Thank You
Sincerely
Louis Gordon-HayIII
1. An instrument store gives a 10% discount to all students off the original cost of an instrument. During a back to school sale an additional 15% is taken off the discounted price. Julie, a student at the local high school, purchases a flute for $306. How much did it originally cost?
A. $325
B. $375
C. $400
D. $408
E. $425
2. If y(x-1)=z then x=
A. y-z
B. z/y + 1
C. y(z-1)
D. z(y-1)
E. 1-zy
3. Which of the following values is NOT equal to 34(58+9)?
A. 34 x 67
B. 58(34+9)
C. 34 x 58 + 34 x 9
D. 1,972 + 306
E. (9 + 58) 34
4. Two angles of a triangle measure 15° and 85 °. What is the measure for the third angle?
A. 50°
B. 55°
C. 60°
D. 80°
E. 90°
5. If 5 ounces is equal to 140 grams, then 2 pounds of ground meat is equal to how many ounces?
A. 863
B. 878
C. 896
D. 915
E. 932
6. Which year did the most children take swimming lessons?

A. 1990
B. 1991
C. 1992
D. 1994
E. 1995
7. Between which year did the largest decrease in children taking swimming lessons occur?
A. 1990-1991
B. 1991-1992
C. 1992-1993
D. 1993-1994
E. 1994-1995
8. What was the average number of children taking swim lessons from 1990 to 1995?
A. 250
B. 308
C. 385
D. 450
E. 1,850
9. Which of the following is equal to 5.93 x 10-2?
A. 0.0593
B. 0.00593
C. 593
D. 5930
E. 59300
10. On a Map, 1 inch represents 20 miles. The distance between 2 towns is 6 1/5 inches. How many miles are actually between the two towns?
A. 65 miles
B. 84 miles
C. 124 miles
D. 138 miles
E. 145 miles
11. Which of the following is a correct graph of x>1, x <4?

A. Line A
B. Line B
C. Line C
D. Line D
E. Line E
12. How many cubed pieces of fudge that are 3 inches on an edge can be packed into a Christmas tin that is 9 inches deep by 12 inches wide by 8 inches high with the lid still being able to be closed?
A. 18
B. 24
C. 32
D. 36
E. 43
13. Sarah is twice as old as her youngest brother. If the difference between their ages is 15 years. How old is her youngest brother?
A. 10
B. 15
C. 20
D. 25
E. 30
14. Which of the following fractions is equal to 5/6?
A. 20/30
B. 15/24
C. 25/30
D. 40/54
E. 2/7
15. What will it cost to tile a kitchen floor that is 12 feet wide by 20 feet long if the tile cost $8.91 per square yard?
A. $224.51
B. $237.60
C. $246.55
D. $271.38
E. $282.32
16. In a writing competition, the first place winner receives ½ of the prize money. The second runner up receives ¼ of what the winner won. What was the total amount of prize money distributed if the winner receives $6,000?
A. $6,000
B. $8,500
C. $12,000
D. $15,000
E. $18,500
17. You are lying 120 ft away from a tree that is 50 feet tall. You look up at the top of the tree. Approximately how far is your hear from the top of the tree in a straight line?
A. 50 feet
B. 75 feet
C. 120 feet
D. 130 feet
E. 150 feet
18. A cyclist bikes x distance at 10 miles per hour and returns over the same path at 8 miles per hour. What is the cyclist's average rate for the round trip in miles per hour?
A. 8.1
B. 8.3
C. 8.6
D. 8.9
E. 9.0
19. If edging cost $2.32 per 12-inch stone, and you want a double layer of edging around your flower bed that is 6 yards by 1 yard. How much will edging you flower bed cost?
A. $32.48
B. $64.96
C. $97.44
D. $129.92
E. $194.88
20. If 3x=6x-15 then x + 8=
A. 5
B. 10
C. 11
D. 12
E. 13
21. The number of milliliters in 1 liter is
A. 10,000
B. 1,000
C. 0.1
D. 0.01
E. 0.001
22. The cost to ride on a ferry is $5.00 per vehicle and driver with an additional cost of 50 cents per passenger. If the charge to get on the ferry is $6.50, how many people were in the vehicle?
A. 1
B. 2
C. 3
D. 4
E. 5
23. What is 1/9 of 9?
A. 1/9
B. 0
C. 1
D. 2
E. 3
24. In his pocket, a boy has 3 red marbles, 4 blue marbles, and 4 green marbles. How many will he have to take out of his pocket to ensure that he has taken out at least one of each color?
A. 3
B. 7
C. 8
D. 9
E. 11
25. Which fraction is equal to 0.20%?
A. 1/20
B. 1/40
C. 1/50
D. 1/400
E. 1/500
26. Find the missing term in the following sequence: 4, 9, 19, __, 79
A. 36
B. 37
C. 38
D. 39
E. 40
27. How much money did Jessica's budget allow for housing in April of 2001?

Answer by nerdybill(7384)

1. An instrument store gives a 10% discount to all students off the original cost of an instrument. During a back to school sale an additional 15% is taken off the discounted price. Julie, a student at the local high school, purchases a flute for $306. How much did it originally cost?
Let x = original cost
from:"gives a 10% discount to all students off the original cost" we get:
.90x
then from:"an additional 15% is taken off the discounted price" we get:
.85(.90x)
and finally from "Julie purchases a flute for $306" we get our equation:
.85(.90x)=306
dividing both sides by .85:
.90x=360
dividing both sides by .90:
.90x=360
x = $400
Ans: C
2. If y(x-1)=z then x=
In this one, simply isolate the variable x to one side of the equation.
y(x-1)=z
first divide both sides by y:
x-1 = z/y
next add 1 to both sides:
x = z/y + 1
Ans: B
3. Which of the following values is NOT equal to 34(58+9)?
This one, you do by elimination...
Ans: B
4. Two angles of a triangle measure 15° and 85 °. What is the measure for the third angle?
The sum of the 3 interior angles of any triangle is 180.
The idea is to add up the two given angles and subtract from 180.
180-(15+85)
=180-(100)
= 80
Ans: D
5. If 5 ounces is equal to 140 grams, then 2 pounds of ground meat is equal to how many grams?
1 pound is 16 ounces
therefore 2 pounds is 32 ounces
32 ounces times 140 grams per 5 ounces:
32(140/5)
= 896 grams
Ans: C
6. Which year did the most children take swimming lessons?
Can't see your graph -- can't answer
7. Between which year did the largest decrease in children taking swimming lessons occur?
Can't see your graph -- can't answer
8. What was the average number of children taking swim lessons from 1990 to 1995?
Can't see your graph -- can't answer
9. Which of the following is equal to 5.93 x 10-2?
10^(-2) is equivalent to .01
therefore
5.93 x .01 = 0.0593

Ans: A
10. On a Map, 1 inch represents 20 miles. The distance between 2 towns is 6 1/5 inches. How many miles are actually between the two towns?
we need to convert 6 1/5 into decimal inches:
1/5 = 0.2 therefore 6 1/5 is 6.2 inches.
Since there is 20 miles per inch:
6.2 x 20 = 124 miles
Ans: C
11. Which of the following is a correct graph of x>1, x <4?
Can't see your figure -- can't answer
12. How many cubed pieces of fudge that are 3 inches on an edge can be packed into a Christmas tin that is 9 inches deep by 12 inches wide by 8 inches high with the lid still being able to be closed?
The volume of the fudge is 3*3*3 = 27 cubic inches
The volume of the pan is 9*12*8 = 864 cubic inches
864/27 = 32
Ans: C
13. Sarah is twice as old as her youngest brother. If the difference between their ages is 15 years. How old is her youngest brother?
Let x = age of youngest brother
then
Sarah age = 2x
from:"difference between their ages is 15 years" we get:
2x - x = 15
x = 15
Ans: B
14. Which of the following fractions is equal to 5/6?
Through elimination -- and because we can multiply numerator and denominator by 5 to get the answer.
Ans: C

15. What will it cost to tile a kitchen floor that is 12 feet wide by 20 feet long if the tile cost $8.91 per square yard?
Convert the 12x20 feet into units of yards:
Area (in yards) = 4(20/3) = 80/3 sq yards
80/3 * 8.91 = $237.60
Ans: B
16. In a writing competition, the first place winner receives ½ of the prize money. The second runner up receives ¼ of what the winner won. What was the total amount of prize money distributed if the winner receives $6,000?
Looks as though this problem gave you MORE information than necessary. From the first sentence alone you can get your answer of 12000.
Ans: C